A rugby player carries the ball 15.9 m straight cl
Answer #1

Here,

Average velocity during the first 3.46 s = 15.9/3.46 = 4.59 m/s
Average velocity during the next 0.293 s = -2.66/0.293 = -9.08 m/s
Average velocity during the final 3.68 s = 17.6/3.68 = 4.78 m/s

Now, the total average velocity = total displacement/ total time
Total displacement = 15.9 – 2.66 + 17.6 = 30.84 m
Total time = 3.46 + 0.293 + 3.68 = 7.433 s
So, overall average velocity = 30.84/7.433 = 4.15 m/s